OpenStudy (anonymous):

Find the equation of the tangent to the curve defined by y =e^x that is perpendicular to the line defined by 3x+y=1.

OpenStudy (kc_kennylau):

There are three pieces of information here: 1. it's a straight line (any tangent must be a straight line) 2. it's tangent to the curve y=e^x 3. it's perpendicular to the line 3x+y=1

OpenStudy (anonymous):

I got y=1/3 e^x but the answer should be x-3y+(1+ln3)=0 How do I get that answer?

OpenStudy (kc_kennylau):

Let's find the slope of the line from the third piece of information

OpenStudy (anonymous):

m=-3 in the line so slope of perpendicular line should be 1/3

OpenStudy (kc_kennylau):

yep

OpenStudy (kc_kennylau):

At which point of the curve y=e^x will the slope of the tangent be 1/3?

OpenStudy (anonymous):

um -1?

OpenStudy (kc_kennylau):

Why so?

OpenStudy (anonymous):

ln1/3?

OpenStudy (kc_kennylau):

I'm asking for a point... A point has 2 coordinates...

OpenStudy (anonymous):

(-1,1)

OpenStudy (kc_kennylau):

where did the ln1/3 go

OpenStudy (kc_kennylau):

this point ain't even on the curve

OpenStudy (anonymous):

lol idk o.o

OpenStudy (kc_kennylau):

At which point of the curve y=e^x will the slope of the tangent be 1/3?

OpenStudy (kc_kennylau):

ln1/3 is near

OpenStudy (anonymous):

I'm not sure.

OpenStudy (anonymous):

3e?

OpenStudy (kc_kennylau):

nope

OpenStudy (kc_kennylau):

never mind

OpenStudy (kc_kennylau):

On the curve y=e^x, when x=ln1/3, y=?

OpenStudy (anonymous):

1/3

OpenStudy (kc_kennylau):

now you can answer

OpenStudy (kc_kennylau):

Now you know that the slope of the line is 1/3, and the line passes through the point (ln1/3,1/3)

OpenStudy (kc_kennylau):

And you should be able to find the equation of the line

OpenStudy (anonymous):

I got x-3y-ln1/3 +1=0

OpenStudy (kc_kennylau):

Mind showing us the steps?

OpenStudy (anonymous):

y=1/3 (x-ln1/3) + 1/3 expanded and multiplied whole thing by 3

OpenStudy (kc_kennylau):

The answers are all equivalent

OpenStudy (anonymous):

How so?